OpenStudy (anonymous):

I NEED HELP QUICKLY!! PLEASE!!! What is the difference of x over the difference of x and 3 − the difference of 5 x and 1 all over the expression x squared minus 6 x plus 9?

OpenStudy (anonymous):

\[\frac{ x }{ x-3 } - \frac{ 5x-1 }{ x^2 - 6x + 9 }\]

OpenStudy (anonymous):

Hey @AriPotta can u help with this?

OpenStudy (anonymous):

@dpaInc ? @AriPotta ? @dumbsearch2 ?

OpenStudy (anonymous):

I need help like ASAP

OpenStudy (mertsj):

\[\frac{x(x-3)}{(x-3)(x-3)}-\frac{5x-1}{(x-3)(x-3)}=\frac{x^2-3x-5x+1}{(x-3)(x-3)}=\frac{x^2-8x+1}{(x-3)(x-3)}\]

OpenStudy (anonymous):

OMG THANK YOU!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (mertsj):

yw